OpenStudy (anonymous):

How to I simplify rational exponents? Such as, (81^1/2 - 9^1/2)^2

OpenStudy (jdoe0001):

\(\Large {\bf \left(81^{\frac{1}{2}}-9^{\frac{1}{2}}\right)^2\implies \left((9^2)^{\frac{1}{2}}-(3^2)^{\frac{1}{2}}\right)^2\\ \quad \\\implies \left(9^{2\cdot \frac{1}{2}}-3^{2\cdot \frac{1}{2}}\right)^2}\)